finding the roots of a quadratic equation with leading coefficent greater than one.
4t+20t=-25
i know that you must first make the equation equal 0. therefore it would b....
4t+20t+25=0
can u please show me in detail what to do next Answer by mananth(16946) 
 4t+20t=-25
Since it is quadratic
4t^2+20t+25=0
4t^2+10t+10t+25=0
2t(2t+5)+5(2t+5)=0
(2t+5)^2=0
2t+5=0
t=-5/2